Suppose that 88% of college men and 84% of college women were employed last summer, based on a sample survey that interviewed SRSs of 400 college men and 400 college women. The two samples were independent.

a) What is the approximate distribution of the proportion p̂F of women who worked last summer? What is the approximate distribution of the proportion p̂M of men who worked?

b) The survey wants to compare men and women. What is the approximate distribution of the difference in the proportions who worked, p̂M - p̂F? EXPLAIN

c) What is the probability that in the sample a higher proportion of women than men worked last summer?
Answer #1

Solution:

An Swea: Given Dala is wonen ten ni - H00 ng - 400 Bi - 0.88 B, 0.54 (a) proposition foor women P2 = 0.84 Standard error - Pu- Plzc-0.04 ) 0.024 = = P(Z < -1.632 ) 0.5 - P (-1.632 < 720) = 0.5 -0.4484 = 0.0516.